# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

### v2

@@ 6.1.2 Singletons
Table 7 shows the results of all proposed variants of the singletons extension. Again, None column represents the baseline xlm-r-large trained from scratch for each dataset. Dummy column shows the results of plain dummy antecedent for singletons. mask extends the previous method binary score masking for singletons. This means that the method does not use the binary score for similarities to the singleton dummy antecedent. separate uses separate FFNN for singletons score prediction (different from the one used for standard mentions). The last method mentions represents the method described in section 4.5.2. Additionally, the Table shows the number of parameters added by each variant of the model. The last column shows the percentage of singleton entities in each dataset (datasets with few singletons do not need any singleton model).
@@
From the table, we can see that on average the best results were achieved with the mention model. Mask model has very poor results, but it makes sense because the number of specific parameters for singletons modeling is very low (3 092 params is used as embedding for dummy antecedent) and if we discard the binary score, we reduce the learning power of the model even more. Dummy model does not help for any dataset statistically significantly. Separate and Mentions models are significantly better for Spanish (Mentions even for Catalan) which has almost no singletons. One reasonable explanation is that the baseline model is under-parametrized for this dataset. We evaluated this hypothesis by increasing the model size, but the improvement was not statistically significant. In the following chapter, we evaluate singletons on the model with joined pretraining, we compute confidence intervals, and in Table 13 we can see that the improvement is below statistical significance (on significance level). So we believe that these strange results in Table 7 are caused by random chance.
@@ 6.1.3 Best Combinations
Table 8 presents the performance of all the extensions proposed. The table also shows the 95% confidence intervals. Span2head and Only head extend Trees (both also uses syntax tree representations). We can see, that tree representations alone improve the results only for ParCor corpora. Span2Head model helps significantly only for English-GUM and Hungarian-korkor corpora but it is not better than only heads model. Singletons modeling helps for most of the datasets that have singletons annotated. The exceptions are in Norwegian Bokmaal and German Potsdam datasets. Another interesting point is that for German ParCor dataset singletons modeling helps significantly even though de-parcor itself does not contain singletons. One possible explanation is in the joined model. Adding the singletons from other languages to the training data can help to predict the mentions in German ParCor dataset. Results in Table 7 support this theory because we can see that when we train the model from scratch, singletons modeling does not help. Only heads model helps for almost all the datasets since it reduces the mention space and improves precision significantly. The extensions selected for the best model are highlighted in green color and the last column shows the results of the best combination. Please note that the results may show slight inconsistencies, even for datasets where a single extension was selected (e.g., ’Only heads’ and ’Best Comb.’ for Hungarian). This is because the experiments were rerun, averaging results across different runs with the same configuration. However, these differences are within the confidence intervals, ensuring the results remain reliable.
Dataset/Extension | Joined | Trees | Span2head | Singletons | Only Heads | Best Comb. |
@@
cs_pcedt | 73.23 .08 | 73.2 .1 | 71.27 .11 | 73.17 .22 | 75.95 .13 | 75.89 .09 |
cs_pdt | 73.47 .23 | 73.2 .46 | 72.98 .29 | 73.37 .28 | 79.43 .1 | 79.51 .02 |
@@
[…]
fr_democrat | 65.24 .49 | 65.46 .4 | 65.38 .49 | 66.23 .44 | 67.72 .81 | 68.58 .23 |
[…]
hu_szegedkoref | 69.52 .59 | 69.59 .54 | 69.32 .34 | 69.47 .82 | 70.68 .46 | 70.67 .31 |
[…]
no_nynorsknarc | 75.98 .17 | 74.79 .48 | 74.96 1 | 76.59 .42 | 78.51 .21 | 78.81 .16 |
[…]
ru_rucor | 70.38 .64 | 70.23 .42 | 70.49 .42 | 70.6 .41 | 75.8 .58 | 75.98 .42 |
tr_itcc | 45.39 1.3 | 34.79 7.7 | 29.53 .5.2 | 45.63 .4 | 45.17 3.9 | 44.17 1.8 |
avg | 70.95 | 70.94 | 69.69 | 71.77 | 73.70 | 74.67 |
6.1.4 Overlapping Segments
Table 9 evaluates longer context implementation through overlapping segments. We use a maximum length of 8 segments666since we use this maximum length in most of our other experiments, and we evaluate minimum and maximum overlap (1 and 7, respectively). We also try variants with filtering already-seen mentions as described in section 4.6 (columns min_filter and max_filter). For most of the datasets, the results do not change at all, so additionally, we run the same experiment with just 4 segments (Table 10)777which is the minimum number of segments we used in our models, and we count basic statistics of long coreference relations (Table 11). Table 11 presents three types of information:
@@ 6.1.4 Overlapping Segments
column nearest coref cross N is the percentage of mentions for which the nearest antecedent is already in a different segments block, so all the antecedents are in the different block and the antecedent cannot be predicted with the standard model.
@@
We can see that a significant number of log-distant coreferences are in French, Turkish, and Russian datasets. Some of them are also in both Czech datasets, Catalan and Polish. By far the most long coreferences are in the French dataset. From Table 9, we can see that segments overlapping decreases the performance. We investigated the dataset closely and found out that in this dataset multiple documents are merged into a single document quite often. When this happens, cross coreferences between those concatenated documents are not annotated so the model which can predict long-distant coreferences is penalized. To evaluate the models correctly, we add the variant where we split the documents in French dev data manually (row fr_split). After splitting, results with segments overlapping improve significantly, but they are still not as high as we expected in comparison with Russian, where the overlap helps significantly. To evaluate this a bit further, we also computed the recall for cross-segment coreference links. The recall is 72% for Russian and only 44% for French. By comparing these datasets manually, we conclude that the coreference chains are very long in the French dataset, and it is probably too hard to predict them correctly from the antecedents (one mistake in the chain might have a huge effect on the overall performance). For the Turkish dataset, it is hard to make any significant improvement due to its low quality.
max. segments 8 | None | min | min_filter | max | max_filter |
[…]
cs_pcedt | 78.23 .13 | 78.18 .11 | 78.23 .15 | 78.19 .15 | 78.12 .1 |
cs_pdt | 80.02 .21 | 79.95 .19 | 79.97 .2 | 79.98 .2 | 79.96 .19 |
de_parcorfull | 81.54 1.3 | 81.54 1.3 | 81.54 1.3 | 81.54 1.3 | 81.54 1.3 |
de_potsdamcc | 75.88 2.2 | 75.88 2.2 | 75.88 2.2 | 75.88 2.2 | 75.88 2.2 |
en_gum | 76.25 .12 | 76.25 .12 | 76.25 .12 | 76.25 .12 | 76.25 .12 |
en_parcorfull | 67.92 3.1 | 67.92 3.1 | 67.92 3.1 | 67.92 3.1 | 67.92 3.1 |
es_ancora | 82.65 .27 | 82.65 .27 | 82.65 .27 | 82.65 .27 | 82.65 .27 |
[…]
hu_korkor | 73.41 .43 | 73.41 .43 | 73.41 .43 | 73.41 .43 | 73.41 .43 |
hu_szegedkoref | 71.2 .14 | 71.2 .14 | 71.2 .14 | 71.2 .14 | 71.2 .14 |
lt_lcc-corefud | 76.82 .41 | 76.82 .41 | 76.82 .41 | 76.82 .41 | 76.82 .41 |
no_bokmaalnarc | 78.96 .37 | 78.96 .37 | 78.96 .37 | 78.96 .37 | 78.96 .37 |
no_nynorsknarc | 80.5 .14 | 80.5 .14 | 80.5 .14 | 80.5 .14 | 80.5 .14 |
[…]
ru_rucor | 75.27 .42 | 77.23 .63 | 76.65 .39 | 76.77 .39 | 77.19 .57 |
[…]
avg | 74.46 | 74.60 | 74.60 | 74.57 | 74.60 |
max. segments 4 | none | min | min filter | max | max filter |
[…]
cs_pcedt | 77.82 .088 | 77.29 .17 | 77.47 .17 | 77.84 .096 | 77.65 .11 |
cs_pdt | 79.9 .22 | 79.87 .16 | 79.84 .2 | 79.91 .22 | 79.74 .19 |
de_parcorfull | 81.54 1.3 | 81.54 1.3 | 81.54 1.3 | 81.54 1.3 | 81.54 1.3 |
de_potsdamcc | 75.88 2.2 | 75.88 2.2 | 75.88 2.2 | 75.88 2.2 | 75.88 2.2 |
en_gum | 76.25 .12 | 76.25 .12 | 76.25 .12 | 76.25 .12 | 76.25 .12 |
en_parcorfull | 67.92 3.1 | 67.92 3.1 | 67.92 3.1 | 67.92 3.1 | 67.92 3.1 |
es_ancora | 82.59 .27 | 82.48 .31 | 82.56 .26 | 82.59 .27 | 82.57 .28 |
fr_democrat | 68.81 .84 | 68.64 .89 | 68.33 .79 | 68.81 .92 | 67.77 1.7 |
[…]
hu_korkor | 73.41 .43 | 73.41 .43 | 73.41 .43 | 73.41 .43 | 73.41 .43 |
hu_szegedkoref | 71.2 .14 | 71.2 .14 | 71.2 .14 | 71.2 .14 | 71.2 .14 |
lt_lcc | 76.82 .41 | 76.82 .41 | 76.82 .41 | 76.82 .41 | 76.82 .41 |
no_bokmaalnarc | 78.96 .36 | 78.91 .38 | 78.87 .34 | 78.96 .36 | 78.47 .37 |
no_nynorsknarc | 79.87 .57 | 80.32 .18 | 80.15 .37 | 79.87 .57 | 78.94 .12 |
pl_pcc | 75.94 .15 | 76.05 .18 | 76.03 .22 | 75.93 .15 | 75.71 .17 |
ru_rucor | 75.58 .37 | 76.54 .31 | 75.89 .55 | 75.62 .42 | 72.8 .25 |
tr_itcc | 43.61 1.5 | 43.85 1.8 | 43.81 1.3 | 43.59 1.5 | 42.19 2.7 |
avg | 74.31 | 74.39 | 74.33 | 74.24 | 73.88 |
4 | 8 | 4 | 8 | 4 | 8 | |||||||
ca_ancora | 5.975 | 1.459 | .6044 | .2747 | 4.739 | 1.422 | ||||||
cs_pcedt | 15.82 | 7.398 | 1.107 | .1441 | 13.55 | 2.133 | ||||||
cs_pdt | 9.787 | 1.54 | .4787 | .06769 | 4.175 | .3976 | ||||||
de_parcorfull | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
de_potsdamcc | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
en_gum | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
en_parcorfull | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
es_ancora | 3.012 | 0.0 | .1396 | 0.0 | .5076 | 0.0 | ||||||
fr_democrat | 52.86 | 25.12 | 1.664 | .8106 | 34.86/27.52 | 27.52/12.84 | ||||||
hu_korkor | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
hu_szegedkoref | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
lt_lcc | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||
no_bokmaalnarc | 9.026 | 0.0 | .1092 | 0.0 | 3.846 | 0.0 | ||||||
no_nynorsknarc | 15.49 | 0.0 | .5405 | 0.0 | 10.29 | 0.0 | ||||||
pl_pcc | 11.79 | 2.724 | .3591 | .1554 | 2.954 | .8439 | ||||||
ru_rucor | 43.77 | 21.76 | 3.273 | 1.283 | 35.44 | 16.46 | ||||||
tr_itcc | 55.6 | 25.07 | 5.703 | 2.037 | 55.56 | 11.11 |
@@ 6.2 Cross-lingual Transfer
Table 12 shows the results of cross-lingual transfer evaluation. We can see that for most of the datasets fine-tuning the model on the specific data helps but the difference is not so significant. For several datasets, the results are even comparable. For small datasets, the fine-tuned models achieve much better results than the joined model. Note that we do not use any weighing of particular datasets in the joined collection so small datasets have a marginal effect on the training. The results of dataset zero-shot and language zero-shot are mostly very predictable. For the languages that have more datasets in CorefUD collection, the drop of zero-shot compared to the joined model was smaller than for the rest of the datasets. When we discard all the datasets for each language, the results are mostly significantly worse compared to discarding just single dataset. This behavior is expected. Several exceptions should be analyzed more. The strangest results are on the Turkish dataset. Here the zero-shot model achieves the best performance. Together with the fact that the numbers are very low for this dataset, it suggests that there is some noise in the dataset annotation. If we train the model on the training part of the Turkish dataset the noise is present on two sides (training and testing). If we train a zero-shot model we eliminate one part.
Another interesting aspect is the superior performance of de-parcor and english-gum dataset in the language zero-shot scenario
@@ 6.3 Final Results
Table 13 presents the final results of the best-performing models on a test set. For comparison, we use the best model of CRAC 2023 Multilingual coreference resolution shared task from [37] but not the best submission. The best model is much larger than ours, and it uses ensembling, so as the main model for comparison, we use their RemBERT version without ensembling, which has a comparable size to our model.
From the table, we can see that our model outperformed CorPipe on most of the datasets. For some datasets (mostly the ones without singletons annotated), it outperformed even the large CorPipe model, which has approximately 3 times more trainable parameters. The only dataset where CorPipe outperformed our model by a large margin is the Polish one, which we consider an anomaly because, for this dataset, CorPipe results on the test set are much better than for the dev set. On the dev set, we achieve similar results. The same happened in the opposite way for the Russian dataset. Surprisingly, we even outperformed the large CorPipe model in average score, but this is caused mainly by a large margin on Turkish and German-parcor.
Our Best | CorPipe | CorPipe-large | |
ca_ancora | 82.05 | 79.93 | 82.39 |
@@
hu_szegedkoref | 69.02 | 65.4 | 69.16 |
lt_lcc | 70.69 | 68.63 | 75.63 |
no_bokmaalnarc | 76.24 | 75.43 | 78.94 |
no_nynorsknarc | 74.54 | 73.64 | 77.24 |

### v3

@@ 6.1.3 Singletons
Table 7 shows the results of all proposed variants of the singletons extension. None column represents the baseline xlm-r-large trained from scratch for each dataset. Dummy column shows the results of plain dummy antecedent for singletons. mask extends the previous method binary score masking for singletons. This means that the method does not use the binary score for similarities to the singleton dummy antecedent. separate uses separate FFNN for singletons score prediction (different from the one used for standard mentions). The last method mentions represents the method described in section 4.5.2. Additionally, the Table shows the number of parameters added by each variant of the model. The last column shows the percentage of singleton entities in each dataset (datasets with few singletons do not need any singleton model).
@@
The table shows that, on average, the best results were achieved with the mention model. Mask model has very poor results. Still, it makes sense because the number of specific parameters for singletons modeling is very low (3 092 params are used as embedding for dummy antecedent). If we discard the binary score, we reduce the learning power of the model even more. Dummy model does not help for any dataset statistically significantly. Separate and Mentions models are significantly better for Spanish (Mentions even for Catalan) which has almost no singletons. One reasonable explanation is that the baseline model is under-parametrized for this dataset. We evaluated this hypothesis by increasing the model size, but the improvement was not statistically significant. In the following chapter, we evaluate singletons on the model with joined pretraining; we compute confidence intervals, and in Table 13, we can see that the improvement is below statistical significance (on significance level). So, we believe that these strange results in Table 7 are caused by random chance.
@@
The reason is probably in a better mention identification due to more training data. We also analyzed an additional metric, mention overlap ratio (MOR), which measures token overlap between gold and system mentions. Singleton modeling increased MOR from 65% to 73% compared to baseline. We investigated concrete examples to validate this hypothesis. An example from the English GUM dataset follows:
- baseline
Research on adult-learned second language (L2e1) has provided considerable insight into the neurocognitive mechanisms underlying the learning and processing of L2e1 grammare2.
- singletons/GOLD
Research on adult-learned second languagee2 (L2e2)e1 has provided considerable insight into the neurocognitive mechanisms underlying the learning and processing of L2e2 grammare6e5e4e3.
We can see that the model with singletons can handle even complicated nested mentions, whereas the baseline cannot.
6.1.4 Headword Representation
We went through some examples on the datasets where heads-only representation helps very significantly. For example, in Czech PDT, we found that there are many very long entity mentions that are not hard to capture with the standard span-based model. Sometimes a mention does not even fit into the maximum span length. We believe this is the main reason why the head-based model performs so much better. For example, in the sentence from the Czech PDT corpus:
"Ústavní soud rozhodl ve středu o podnětu skupiny 19 členů pražských zastupitelstev na vypuštění té části zákona o volbách do obecních zastupitelstev, která podmiňuje registraci sdružení nezávislých kandidátů předložením petice podepsané sedmi procenty obyvatel obce."
In English:
"On Wednesday the Constitutional Court ruled on a petition filed by a group of 19 members of the Prague City Council to strike out the part of the law on municipal elections that makes the registration of an association of independent candidates contingent on the submission of a petition signed by seven percent of the municipality’s residents."
The whole sentence is a single entity representing the concrete decision of the court, and there are many mentions of this decision later in the document.
@@ 6.1.5 Best Combinations
Table 8 presents the performance of all the extensions proposed. The table also shows the 95% confidence intervals. Span2head and Only head extend Trees (both also uses syntax tree representations). We can see, that tree representations alone improve the results only for ParCor corpora. Span2Head model helps significantly only for English-GUM and Hungarian-korkor corpora but it is not better than only heads model. Singletons modeling helps for most of the datasets that have singletons annotated. The exceptions are in the Norwegian Bokmaal and German Potsdam datasets. Another interesting point is that for the German ParCor dataset, singletons modeling helps significantly even though de-parcor itself does not contain singletons. One possible explanation is in the joined model. Adding the singletons from other languages to the training data can help to predict the mentions in the German ParCor dataset. Results in Table 7 support this theory because we can see that when we train the model from scratch, singletons modeling does not help. Only heads model helps for almost all the datasets since it reduces the mention space and improves precision significantly. The extensions selected for the best model are highlighted in green color and the last column shows the results of the best combination. Please note that the results may show slight inconsistencies, even for datasets where a single extension was selected (e.g., ’Only heads’ and ’Best Comb.’ for Hungarian). This is because the experiments were rerun, averaging results across different runs with the same configuration. However, these differences are within the confidence intervals, ensuring the results remain reliable.
6.1.6 Overlapping Segments
@@
cross N segment corefs [%] nearest coref cross N [%] segments over N [%] 4 8 4 8 4 8 ca_ancora 5.975 1.459 .6044 .2747 4.739 1.422 cs_pcedt 15.82 7.398 1.107 .1441 13.55 2.133 cs_pdt 9.787 1.54 .4787 .06769 4.175 .3976 de_parcorfull 0.0 0.0 0.0 0.0 0.0 0.0 de_potsdamcc 0.0 0.0 0.0 0.0 0.0 0.0 en_gum 0.0 0.0 0.0 0.0 0.0 0.0 en_parcorfull 0.0 0.0 0.0 0.0 0.0 0.0 es_ancora 3.012 0.0 .1396 0.0 .5076 0.0 fr_democrat 52.86 25.12 1.664 .8106 34.86/27.52 27.52/12.84 hu_korkor 0.0 0.0 0.0 0.0 0.0 0.0 hu_szegedkoref 0.0 0.0 0.0 0.0 0.0 0.0 lt_lcc 0.0 0.0 0.0 0.0 0.0 0.0 no_bokmaalnarc 9.026 0.0 .1092 0.0 3.846 0.0 no_nynorsknarc 15.49 0.0 .5405 0.0 10.29 0.0 pl_pcc 11.79 2.724 .3591 .1554 2.954 .8439 ru_rucor 43.77 21.76 3.273 1.283 35.44 16.46 tr_itcc 55.6 25.07 5.703 2.037 55.56 11.11
Table 9 evaluates longer context implementation through overlapping segments. We use a maximum length of 8 segments777since we use this maximum length in most of our other experiments, and we evaluate minimum and maximum overlap (1 and 7, respectively). We also try variants with filtering already-seen mentions as described in section 4.6 (columns min_filter and max_filter). For most of the datasets, the results do not change at all, so additionally, we run the same experiment with just 4 segments (Table 10)888which is the minimum number of segments we used in our models, and we count basic statistics of long coreference relations (Table 11). Table 11 presents three types of information:
@@
Column nearest coref cross N is the percentage of mentions for which the nearest antecedent is already in a different segments block, so all the antecedents are in the different block, and the antecedent cannot be predicted with the standard model.
@@
A significant number of log-distant coreferences are in French, Turkish, and Russian datasets. Some of them are also in both Czech datasets, Catalan and Polish. By far the most long coreferences are in the French dataset. From Table 9, we can see that segments overlapping decreases the performance. We investigated the dataset closely and found out that in this dataset, multiple documents are merged into a single document quite often. When this happens, cross coreferences between those concatenated documents are not annotated so the model which can predict long-distant coreferences is penalized. To evaluate the models correctly, we add the variant where we split the documents in French dev data manually (row fr_split). After splitting, results with segments overlapping improve significantly, but they are still not as high as we expected in comparison with Russian, where the overlap helps significantly. To evaluate this a bit further, we also computed the recall for cross-segment coreference links. The recall is 72% for Russian and only 44% for French. By comparing these datasets manually, we conclude that the coreference chains are very long in the French dataset, and it is probably too hard to predict them correctly from the antecedents (one mistake in the chain might have a huge effect on the overall performance). For the Turkish dataset, it is hard to make any significant improvement due to its low quality.
@@ 6.2 Cross-lingual Transfer
Table 12 shows the results of cross-lingual transfer evaluation. We can see that for most of the datasets, fine-tuning the model on the specific data helps, but the difference is not so significant. For several datasets, the results are even comparable. For small datasets, the fine-tuned models achieve much better results than the joined model. Note that we do not use any weighing of particular datasets in the joined collection, so small datasets have a marginal effect on the training. The results of dataset zero-shot and language zero-shot are mostly very predictable. For the languages with more datasets in CorefUD collection, the drop of zero-shot compared to the joined model was smaller than for the rest of the datasets. When we discard all the datasets for each language, the results are mostly significantly worse compared to discarding just a single dataset. This behavior is expected. Several exceptions should be analyzed more. The strangest results are on the Turkish dataset. Here, the zero-shot model achieves the best performance. Together with the fact that the numbers are very low for this dataset, it suggests that there is some noise in the dataset annotation. If we train the model on the training part of the Turkish dataset the noise is present on two sides (training and testing). If we train a zero-shot model, we eliminate one part.
Another interesting aspect is the superior performance of de-parcor and english-gum datasets in the language zero-shot scenario.
@@ 6.3 Final Results
| Our Best | CorPipe | CorPipe-large | wl-coref | |
| ca_ancora | 82.05 | 79.93 | 82.39 | - |
@@
| hu_szegedkoref | 69.02 | 65.4 | 69.16 | - |
| lt_lcc | 70.69 | 68.63 | 75.63 | - |
| no_bokmaalnarc | 76.24 | 75.43 | 78.94 | - |
| no_nynorsknarc | 74.54 | 73.64 | 77.24 | - |
@@
| OntoNotes | 80.5 | 80.7 | 82 | 81 |
Table 13 presents the final results of the best-performing models on a test set. For comparison, we use the best model of CRAC 2023 Multilingual coreference resolution shared task from [31] but not the best submission. The best model is much larger than ours, and it uses ensembling, so as the main model for comparison, we use their RemBERT version without ensembling, which has a comparable size to the proposed model.
We can see that the proposed model outperformed CorPipe on most of the datasets. For some datasets (mostly the ones without singletons annotated), it outperformed even the large CorPipe model, which has approximately 3 times more trainable parameters. The only dataset where CorPipe outperformed our model by a large margin is the Polish one, which we consider an anomaly because, for this dataset, CorPipe results on the test set are much better than for the dev set. On the dev set, we achieve similar results. The same happened in the opposite way for the Russian dataset. Surprisingly, we even outperformed the large CorPipe model in average score, but this is caused mainly by a large margin on Turkish and German-parcor.
We also evaluate our model on English OntoNotes dataset [32] to validate that our model achieves reasonable performance. It should achieve similar results to word-level coreference since other improvements (like singletons modeling) are not important for OntoNotes. We can see that even though our model is slightly worse than CorPipe, the results are reasonable and close enough to the word-level coreference. We believe that the superior performance of wl-coref can be caused by using specific features for OntoNotes in wl-coref (like speaker information). Note that our goal was not to optimize the results on OntoNotes, so we did not optimize the model specifically for this corpus.